OpenStudy (anonymous):

Factor completely. xy-2x+4y-8

OpenStudy (anonymous):

(x+4)(y-2)

OpenStudy (anonymous):

xy-2x+4y-8 =x(y-2)+4(y-2) =(x+4)(y-2)

OpenStudy (anonymous):

xy+2x+4y+8?

OpenStudy (anonymous):

yamaka just answered it... thrice. Factor by grouping.